# ShanghaiTech Mapping Robot is All You Need: Robot System for Collecting Universal Ground Vehicle Datasets

### v2

@@ IV-B Computation Cluster
To control all the sensors and record the data captured by them, a powerful computer is essential for our mapping robot. [5] Since there are too many sensors on the mapping robot while each of them could capture massive amount of data, even the most powerful portable computer does not have enough I/O and storage bandwidth to handle the recording of data from all sensors. A 16-node computation cluster is specially designed for this usage.
The cluster consists of 16 computing nodes, each modified from an ASUS Mini PN51 computer, having a 8-core, 16-threads Ryzen 7 5700U processor, 32 GiB of RAM, 2 TiB of NVME M.2 SSD storage (Samsung 970EVO Plus, sequential write speed up to 3,300 MB/s), 2 Gigabit Ethernet cards, 3 USB 3.2 Gen1 Type-A ports and 2 USB 3.2 Gen2 Type-C ports. The nodes are stacked in a customized case which also accommodates several DC-DC converters and a LAN switch providing power and interconnection for the cluster, as showed in Fig. 4.
@@
Among the nodes there is one master node that acts as the controller of the whole cluster while the other worker nodes do not have their operating system install locally but utilize network boot and use the system image hosted on the master node, making it much easier to configure the sensor drivers and orchestration programs over all 16 nodes. This approach also saves storage on worker nodes, allowing a longer record time.
Each worker node is intended to control and record up to 4 sensors, including 3 through USB 3.0 connection and one through Gigabit Ethernet connection. Sensor drivers help configure the sensors connected to the according node and collect the captured data. Raw sensor data will be saved in situ to the SSD mounted on the worker node and can, on demand, be streamed to the master node through ROS. Recorded sensor data can be collected for post-processing by either copying them to external hard drives connected to worker nodes or forwarding them through the master node to an external server.
@@ IV-C Synchronization Board
To eliminate or remedy the time difference of data collected by different sensors, either electronic synchronization or IEEE 1588 PTP (Precision Time Protocol) is the most preferred approach. Both of them can achieve sub-microsecond precision. We employed Syncboard, a combination of Raspberry Pi 4 computer and specially designed I/O shield, to act as a master clock of the whole robot that provides triggering or time base for sensors equipped as well as computation nodes in the cluster. On the Syncboard there are 12 triggering output channels and a LiDAR synchronization channel, capable of connecting at most 130 sensors.
@@
Each of the 12 triggering output channels can generate square wave signals according to the channel-wise designated parameters including frequency (up to 1000Hz), triggering edge polarity (rising edge, failing edge or both edges), time offset and duty ratio, adapting to various triggering-compatible sensors like cameras and IMU. Voltage level of each channel can be set to 3.3V or 5V by changing the soldering bridge on the board. One channel may have 6 to 40 vacancies available for sensor connection, enabling us to easily trigger multiple sensors with the same signal.
The LiDAR synchronization channel, on the other hand, is capable of generating GPRMC+PPS signal. GPRMC, or the Recommended Minimum Specific GPS/Transit Data, is an NMEA 0183 message type designed for carrying positioning data, including a timestamp accurate to the second. It is often given along with a PPS (Pulse-per-Second) signal that indicates the exact beginning of the second contained in the message. Conventional LiDARs consume GPRMC+PPS signal, often from a real GPS receiver in old days, to set their internal clock. The channel can be configured to have certain baud rate as well as an original or inverted voltage level to adapt to different LiDARs. The channel provides 10 vacencies for LiDAR connection.
Nevertheless, Syncboard broadcasts its clock using PTP throughout the cluster LAN such that all computing nodes can synchronize themselves as well as the sensors connected to them. Syncboard itself takes either GNSS or public NTP server as its external master clock.
Syncboard also provides a handy LED button for us to physically start or stop triggering and check its working status, no need to send commands on remote devices.
@@ IV-D Power Box
Figure 6: The power box and GNSS receiver of ShanghaiTech Mapping Robot
Inside the power box, multiple DC-DC converters turn the battery power rail into 4 stable DC rails, 5V, 12V, 19V, 24V respectively. The 4 converted rails along side with the battery power rail are connected to the power distribution board, showed in Fig. 7, at the top of the power box. Sensors get electricity from sockets on this board unless they use a single USB port for both data and powering. In that case, they get power from cluster nodes. The UGV base and the computation cluster have their power line connected to the back of the power box to get battery power rail directly.
@@ V System Integration & Sensor Setup
With all sensors and backbone hardware ready, in this section we will introduce how we integrated all components together as well as the setup of each sensor. As showed in Fig. 8, we designed a sensor deck, a front sensor tower and a back sensor tower to accommodate all sensors
@@ V-A Front Sensor Tower
The front sensor tower consists of multiple floors, each of them is made of either aluminum or plastic boards.
On the top there mounts the front GNSS antenna as well as 2 rotational LiDARs, Hesai PandarQT and the first (front) Robosense Ruby.
As showed in Fig. 10 and Fig. 15, PandarQT deserves a position on the very top to observe robot’s near surrounding, including objects at very low angle, with no obstruction. Stayed straight under PandarQT, the front Ruby covers far away objects. Both LiDARs, as well as all rest of them, are connected to separate cluster nodes through Ethernet cable. Each of these nodes handles the control and data acquisition of the respective LiDAR. Robosense LiDARs are synced using the LiDAR synchronization channel mentioned in Sec. IV-C while LiDARs of other models use PTP protocol to get time base from the syncboard relayed by the cluster nodes.
One floor lower, the first (front) Robosense Bpearl is installed on the front panel. Its zenith of the FoV points forward, providing a hemispherical view of the front of the robot, also showed in Fig. 15.
Behind the front Bpearl, 2 Grasshopper3 cameras, one Azure Kinect DK and a mmWave Radar, showed in Fig. 10 and Fig. 9 are attached on the two side panels and the back panel of this floor. They look upwards and provide a view of ceilings or treetop above. Inside this floor, there stashes the cables and I/O boxes of all LiDARs on the robot.
@@
These 2 up-looking Grasshopper3 cameras, as well as their front-, left-, right- and back-looking counterparts, are equipped with Kowa LM6JC wide-angle lenses. It has a focal length of 6mm, providing a wide FoV of $81.9^{\circ}\times 61.2^{\circ}$ and a deep depth of view. [41] We set the focal distance to $\infty$ and aperture to their maximum value of $f/1.4$ to ensure a sharp imaging on far surrounding objects and let sufficient amount of light reach the sensors, enabling high frame-rate capturing. These 10 cameras run in hardware triggering mode and are connected to the same triggering channel on the Syncboard, thus capture images at the same time, with same frame rate.
@@
Another floor lower, there are 2 EvC3A event-based cameras, MTi-630R IMU, Livox Avia LiDAR, and 13 KS103 ultrasound range finders. The 2 EvC3A event-based cameras form a stereo pair looking forward, using same lenses as Grasshopper3’s. The focal distance setting is of the same but the aperture is set to the smallest since event-based sensors capture the change of brightness instead of the absolute brightness and work better with reasonably limited lighting. The 2 event-cameras synchronize their internal clock with each other using a 1 MHz signal and also receive triggering signals from Syncboard. However, instead of following triggering signals to capture whole frames as frame-based cameras would do, event-based cameras percept them as ’External Trigger’ events which would be stored in the same stream along with ’Contrast Detector’ events mentioned in Sec. III-D.
MTi-630R is configured to measure linear acceleration, angular velocity, orientation as well as magnetic heading, atmospheric pressure and temperature, all at a rate of up to 400Hz triggered by Syncboard. KS103 ultrasound range finders are installed on curved racks attached to the tower to achieve evenly angled detecting directions. All 13 of them are connected via an I2C bus to a microcontroller which triggers them in an interleaved sequence to avoid cross-talk.
At the bottom of front sensor tower, 2 A315 IR cameras and the Neuvition M1 LiDAR are mounted on the front end of the sensor deck. A315 is equipped with a 18mm lens providing a 25° FoV and has motorized focus motor enabling automatic focusing control. We configured them to output a 16-bit linear temperature image at a frame rate of 60 Hz. They are connected over a Gigabit Ethernet connection to designated cluster nodes that handle their recording and controlling.
At the front of the UGV base, there is a downward looking Grasshopper3 camera, as showed in Fig. 11. Its focal distance is set to match the distance to the ground in order to get a clear view of the floor, road, etc. Alongside it, two 1.5-watt LED lights are installed to ensure enough lighting, enabling a even higher frame rate to compensate for the limited FoV (about $40\times 60$ cm on the ground) due to its low mounting position. We dedicated a triggering channel of Syncboard for this camera to drive it at a different, often faster, frame rate than its counterparts looking around and upward.
@@ V-B Sensor Deck
The sensor deck is made out of one aluminum plate and forms the length and width of the vehicle. It’s mounted directly onto the base UGV by being bolted to the main aluminum extrusion frame around the top of UGV’s payload bay and further supported by aluminum pillars attached to the front and rear bumper bars. Both the front and back sensor tower are attached to the sensor deck. The computation cluster, power box and one out of the 2 batteries are placed through a slot of sensor deck, marked yellow in Fig. 12, onto the UGV’s payload bay.
@@
The sensor deck accommodates the Grasshopper3 cameras, Azure Kinect DK depth cameras and the mmWave radars facing the rest 4 (front, back, left and right) directions. In each direction, the 2 separate Grasshopper3 mounting slots, marked blue in Fig. 12, are capable of aligning the direction and sensor plane of cameras, forming them into a stereo pair.
As showed in Fig. 13, each mmWave radar is mounted on the back of a Azure Kinect DK using a 3D printed part, forming a Kinect-mmWave pair. Such pairs are attached to Kinect mounting slots of sensor deck, marked green in Fig. 12. Such approach insures a uniform translation between all 5 pairs thus helps simplifying the calibration of mmWave radar extrinsics. While mmWave radars do not support hardware triggering, Azure Kinect DKs do. They are all configured to run in the subordinate mode and have their sync-in ports connected to a dedicated triggering channel of Syncboard generating VSync signal at desired frame rate.
@@
The real-time kinematics GNSS receiver is placed on the sensor deck near the back sensor tower. Its 2 antennas are installed atop the front and back sensor tower respectively. The receiver is connected to the cluster LAN and thus has access to online differential corrections data via a 5G cellular link from either local base station or third-party CORS (Continuously Operating Reference Stations) network. It also sends PPS signal to Syncboard via a dedicated cable when acts as an external master clock as mentioned in Sec. IV-C.
The powering, data and synchronization cables of the sensors are routed on the bottom side of the deck.
@@ V-C Back Sensor Tower
The back sensor tower is made out of a single 40mm*40mm aluminum extrusion pillar. Along the pillar there attaches 2 rotational LiDARs, the omnidirectional camera, the rear GNSS antenna as well as WiFi, cellular and joystick receivers.
@@
The LadyBug5+ is mounted at the top of the back sensor tower. The built-in lenses of LadyBug5+ have a $f/2.5$ aperture and a 4.4mm focal length, achieving a FoV of 90% full sphere (6 sensors combined) and acceptable sharpness from about 60cm to infinity.[35] We capture original pinhole images from the 6 sensors rather than using its SDK to stitched panorama images on the fly. It is triggered via a dedicated channel of Syncboard. The back GNSS antenna is placed on the top of LadyBug5+ in a particular orientation that hides itself in a blind spot.
At the middle of the back sensor tower there mount the second (back) Robosense Ruby and Bpearl. are installed on the robot. The back Bpearl simply points the zenith of its FoV backwards, opposite to its front counterpart, while the back Ruby is installed vertically, pitched up 90°, achieving a dense vertical scanning and coverage of far objects with high elevation angle like tall buildings along the routes. In Fig. 15, we show the FoV of all 7 LiDARs on the robot.
@@
A USB hub is installed on the back sensor tower to accommodate WiFi, cellular and joystick receivers, providing connectivity to the robot for controlling, monitoring, and access to online reference servers for time and differential GNSS.
Under the back sensor tower, a second battery is installed in the back bay of UGV.
With everything connected, the communication and synchronization architecture of the whole robot are as shown in Fig. 16 and Fig. 17

### v3

@@ IV-B Computation Cluster
To control all the sensors and record the data captured by them, a powerful computer is essential for our mapping robot. Since there are too many sensors on the mapping robot, each of them capturing massive amounts of data, even the most powerful portable computer does not have enough I/O and storage bandwidth to handle the recording of the data from all the sensors. Thus a very compact, 16-node computation cluster was specially designed for this usage, whose development has been described in [5].
The cluster consists of 16 computing nodes, each modified from an ASUS Mini PN51 computer, having an 8-core, 16-threads Ryzen 7 5700U processor, 32 GiB of RAM, 2 TiB of NVME M.2 SSD storage (Samsung 970EVO Plus, sequential write speed up to 3,300 MB/s), 2 Gigabit Ethernet ports (one on the main board, one in an M2 slot, replacing the WiFi), 3 USB 3.2 Gen1 Type-A ports and 2 USB 3.2 Gen2 Type-C ports. The nodes are stacked in a customized case which also accommodates several DC-DC converters and a LAN switch providing power and interconnection for the cluster, as shown in Fig. 4.
@@
Among the nodes there is one master node that acts as the controller of the whole cluster, while the other worker nodes do not have their operating system installed locally but utilize network boot and use the system image hosted on the master node, making it much easier to configure the sensor drivers and orchestration programs over all 16 nodes. This approach also saves storage on worker nodes, allowing a longer record time.
Each worker node is intended to control and record up to 4 sensors, including 3 through USB 3.0 connections and one through a Gigabit Ethernet connection. Sensor drivers help configure the sensors connected to the according node and collect the captured data. Raw sensor data will be saved in situ to the SSD mounted on the worker node and can, on demand, be streamed to the master node through ROS. Recorded sensor data can be collected for post-processing by either copying them to external hard drives connected to worker nodes or forwarding them via Gigabit Ethernet through the master node to an external server.
@@ IV-C Synchronization Board
To eliminate or remedy the time difference of data collected by different sensors, both electronic synchronization and IEEE 1588 PTP (Precision Time Protocol) are the most preferred approaches. Both of them can achieve sub-microsecond precision. We developed Syncboard, a combination of a Raspberry Pi 4 computer and a specially designed I/O shield, to act as a master clock of the whole robot that provides triggering or the time base for the sensors equipped as well as the computation nodes in the cluster. On the Syncboard there are 12 triggering output channels and a LiDAR synchronization channel, capable of connecting at most 130 sensors.
@@
Each of the 12 triggering output channels can generate square wave signals according to the channel-wise designated parameters including frequency (up to 1000Hz), triggering edge polarity (rising edge, failing edge or both edges), time offset and duty ratio, adapting to various triggering-compatible sensors like cameras and IMU. The voltage level of each channel can be set to 3.3V or 5V by changing the soldering bridge on the board. One channel may have 6 to 40 vacancies available for sensor connection, enabling us to easily trigger multiple sensors with the same signal.
The LiDAR synchronization channel, on the other hand, is capable of generating GPRMC+PPS signal. GPRMC, or the Recommended Minimum Specific GPS/Transit Data, is an NMEA 0183 message type designed for carrying positioning data, including a timestamp accurate to the second. It is often given along with a PPS (Pulse-per-Second) signal that indicates the exact beginning of the second contained in the message. Conventional LiDARs consume the GPRMC+PPS signal, often from a real GPS receiver in old days, to set their internal clock. The channel can be configured to have a certain baud rate as well as an original or inverted voltage level to adapt to different LiDARs. The channel provides 10 vacancies for LiDAR connection.
Additionally, Syncboard broadcasts its clock using PTP throughout the cluster LAN such that all computing nodes can synchronize themselves, as well as the sensors connected to them. Syncboard itself takes either GNSS or public NTP server as its external master clock.
Syncboard also provides a handy LED button for us to physically start or stop triggering and check its working status, no need to send commands on remote devices. Syncboard itself is integrated with the computation cluster. Its I/O shield acts as a part of cluster’s top I/O plate and the Raspberry Pi is accommodated inside the cluster and connected to the Ethernet switch as the 17th node.
@@ IV-D Power Box
Figure 6: The power box and GNSS receiver of ShanghaiTech Mapping Robot.
Inside the power box, multiple DC-DC converters turn the battery power rail into 4 stable DC rails, 5V, 12V, 19V, 24V respectively. The 4 converted rails along side with the battery power rail are connected to the power distribution board, shown in Fig. 7, at the top of the power box. Sensors get electricity from sockets on this board unless they use a single USB port for both data and powering. In that case, they get power from cluster nodes. The UGV base and the computation cluster have their power line connected to the back of the power box to get battery power rail directly.
@@ V System Integration & Sensor Setup
With all sensors and backbone hardware ready, in this section we will introduce how we integrated all components together as well as the setup of each sensor.
The end goal for the robot is to collect very accurate sensor data. For this we will calibrate the extrinsics (the poses) of all sensors, which is described in Section VI. To facilitate this, the sensors need to be rigidly mounted in such a way, that they do not move even when the base is driving over harsh terrain. We thus mechanically designed a very stiff mounting structure. As shown in Fig. 8, the robot features a sensor deck, a front sensor tower and a back sensor tower to accommodate all sensors. The back tower is a single 80 cm tall beam. This beam is screwed into the sensor deck, also with triangle brackets, secured by a back plate to minimize roll rotation and further stiffened by a diagonal aluminum profile connected to the sensor deck as well as a horizontal aluminium profile connected to the front tower to reduce pitch rotation. Fig. 6 shows these support structures. The 60 cm tall front tower is a four-legged aluminum profile structure.
@@ V-A Front Sensor Tower
The front sensor tower consists of multiple floors, each of them is made of either aluminum or acrylic plastic boards.
On the top mounts the front GNSS antenna as well as 2 rotational LiDARs, Hesai PandarQT and the first (front) Robosense Ruby.
As shown in Fig. 10 and Fig. 15, PandarQT deserves a position on the very top to observe robot’s near surrounding, including objects at very low angle, with no obstruction. Located straight under the PandarQT, the front Ruby covers far away objects. Both LiDARs, as well as all rest of them, are connected to separate cluster nodes through Ethernet cables. Each of these nodes handles the control and data acquisition of the respective LiDAR. Robosense LiDARs are synced using the LiDAR synchronization channel mentioned in Sec. IV-C, while LiDARs of other models use PTP protocol to get time base from the syncboard relayed by the cluster nodes.
One floor lower, the first (front) Robosense Bpearl is installed on the front panel. Its zenith of the field of view (FoV) points forward, providing a hemispherical view of the front of the robot, also shown in Fig. 15.
Behind the front Robosense Bpearl, 2 Grasshopper3 cameras, one Azure Kinect DK and a mmWave Radar, shown in Fig. 10 and Fig. 9 are attached on the two side panels and the back panel of this floor. They look upwards and provide a view of ceilings or treetop above. Inside this floor, there stashes the cables and I/O boxes of all LiDARs on the robot.
@@
The two up-looking Grasshopper3 cameras, as well as their front-, left-, right- and back-looking counterparts, are equipped with Kowa LM6JC wide-angle lenses. They have a focal length of 6mm, providing a wide FoV of $81.9^{\circ}\times 61.2^{\circ}$ and a deep depth of view [43]. We set the focal distance to $\infty$ and the aperture to their maximum value of $f/1.4$ to ensure a sharp imaging on far surrounding objects and let sufficient amount of light reach the sensors, enabling high frame-rate capturing. These 10 cameras run in hardware triggering mode and are connected to the same triggering channel on the Syncboard, thus capture images at the same time, with same frame rate.
@@
Yet another floor lower, there are two EvC3A event-based cameras, an MTi-630R IMU, Livox Avia LiDAR, and 13 KS103 ultrasound range finders. The two EvC3A event-based cameras form a stereo pair looking forward, using same wide FoV lenses as Grasshopper3’s. The focal distance setting is the same but the aperture is set to the smallest since event-based sensors capture the change of brightness instead of the absolute brightness and work better with reasonably limited lighting. The two event-cameras synchronize their internal clock with each other using a 1 MHz signal and also receive pulse signals for time-stamping from the Syncboard. Instead of following the signals to capture whole frames as frame-based cameras would do, event-based cameras perceive them as ’External Trigger’ events which would be stored in the same stream along with ’Contrast Detector’ events mentioned in Sec. III-D.
The MTi-630R IMU is configured to measure linear acceleration, angular velocity, orientation and magnetic heading in all 3 axis, as well as atmospheric pressure and temperature, all at a rate of up to 400Hz triggered by the Syncboard. The KS103 ultrasound range finders are installed on curved brackets attached to the tower to achieve evenly angled detecting directions. All 13 of them are connected via an I2C bus to a microcontroller which triggers them in an interleaved sequence to avoid cross-talk.
At the bottom of front sensor tower, two A315 IR cameras and the Neuvition M1 LiDAR are mounted on the front end of the sensor deck. The A315 are equipped with 18mm lenses providing a 25° FoV and have motorized focus motors enabling automatic focusing control. We configured them to output a 16-bit linear temperature image at a frame rate of 60 Hz. They are connected over a Gigabit Ethernet connection to designated cluster nodes that handle their recording and controlling.
At the front of the UGV base, there is a downward looking Grasshopper3 camera, as shown in Fig. 11. Its focal distance is set to match the distance to the ground in order to get a clear view of the floor, road, etc. Alongside it, two 1.5-watt LED lights are installed to ensure enough lighting, enabling an even higher frame rate to compensate for the limited FoV (about $40\times 60$ cm on the ground) due to its low mounting position. We dedicated a triggering channel of the Syncboard for this camera to drive it at a different, often faster, frame rate than its counterparts looking around and upward.
@@ V-B Sensor Deck
The 94cm x 63cm sensor deck is CNCed from aircraft grade 7050 aluminum of 10mm thickness and anodized to reduce reflections that may interfere with the sensor data. The deck’s CAD is available alongside the dataset. This deck defines the length and width of the vehicle. It it mounted directly onto the base UGV by being bolted to the main aluminum extrusion frame around the top of UGV’s payload bay by two 20mm x 80mm x 380mm aluminum profiles and further supported by aluminum pillars attached along with the front and rear bumper bars to the main chassis of the Husky UGV. Both the front and back sensor towers are attached to the sensor deck. The computation cluster, power box and one of the two batteries are placed through a 45cm x 38cm opening of the sensor deck, marked yellow in Fig. 12, into the UGV’s payload bay.
@@
The sensor deck accommodates the Grasshopper3 cameras, the Azure Kinect DK depth cameras and the mmWave radars facing the front, back, left and right directions. In each direction, the two separate Grasshopper3 mounting slots, marked blue in Fig. 12, are capable of aligning the direction and sensor plane of cameras, forming them into a stereo pair.
As shown in Fig. 13, each mmWave radar is mounted on the back of a Azure Kinect DK using a 3D printed part, forming a Kinect-mmWave pair. Such pairs are attached to Kinect mounting slots of sensor deck, marked green in Fig. 12. Such approach ensures a uniform translation between all 5 pairs and thus helps simplifying the calibration of mmWave radar extrinsics. While mmWave radars do not support hardware triggering, Azure Kinect DKs do. They are all configured to run in the subordinate mode and have their sync-in ports connected to a dedicated triggering channel of the Syncboard generating the Kinect’s VSync signal at desired frame rates.
@@
The real-time kinematics GNSS receiver is placed on the sensor deck near the back sensor tower. Its two antennas are installed atop the front and back sensor tower, respectively. The receiver is connected to the cluster LAN and thus has access to online differential corrections data via a 5G cellular link from either local base station or the third-party CORS (Continuously Operating Reference Stations) network. It also sends a PPS signal to the Syncboard via a dedicated cable, such that it can act as an external master clock, if we configure it this way, as mentioned in Sec. IV-C.
The powering, data and synchronization cables of the deck’s sensors are routed in cable channels on the bottom sides of the deck.
@@ V-C Back Sensor Tower
The back sensor tower is made out of a single $40\times 40\times 800$ mm aluminum profile pillar. Along the pillar, two rotational LiDARs, the omni-directional camera, the rear GNSS antenna as well as WiFi, cellular and joystick receivers are attached.
@@
The omni-directional LadyBug5+ is mounted at the top of the back sensor tower. The built-in lenses of LadyBug5+ have a $f/2.5$ aperture and a 4.4mm focal length, achieving a FoV of 90% full sphere (6 sensors combined) and acceptable sharpness from about 60cm to infinity [35]. We capture original pinhole images from the 6 sensors rather than using its SDK to stitch panorama images on the fly, for preserving the raw data and to save computation time. It is triggered via a dedicated channel of Syncboard. The back GNSS antenna is placed on the top of LadyBug5+ in a particular orientation that hides it in a blind spot.
The second back Robosense Ruby and Bpearl are mounted in the middle of the tower. The back Bpearl simply points the zenith of its FoV backwards, opposite to its front counterpart, while the back Ruby is installed vertically, pitched up 90°, achieving a dense vertical scanning and coverage of far objects with high elevation angle like tall buildings along the routes. In Fig. 15, we show the FoV of all 7 LiDARs on the robot.
@@
A USB hub is installed on the back sensor tower to accommodate WiFi, cellular and joystick receivers, providing connectivity to the robot for controlling, monitoring, and access to online reference servers for time and differential GNSS. Below the back sensor tower, a second battery is installed in the back bay of UGV. With everything connected, the communication and synchronization architectures of the whole robot are as shown in Fig. 16 and Fig. 17.